OpenStudy (anonymous):

Find the common solution by graphing. 2x + y> 3 x< 1

OpenStudy (anonymous):

y=3x 2. x=-y 2x+y=-2 3. 2x+2y=4 x=10-3y 4. y=7-x 2x+3y=-1